OpenStudy (anonymous):

What is the solution to the equation 3 1 – 2x = 2 ? x ≈ –0.292 x ≈ –0.185 x ≈ 0.185 x ≈ 0.292

OpenStudy (ghazi):

31-2=2x...now solve for x

OpenStudy (ghazi):

is it 3(1-2x)=2

jimthompson5910 (jim_thompson5910):

Is the equation \[\Large 3^{1-2x} = 2\] ???

OpenStudy (anonymous):

You must have written the problem down wrong.

OpenStudy (anonymous):

no

OpenStudy (ghazi):

can you rewrite your equation?

OpenStudy (anonymous):

What is the solution to the equation 3 ^1 – 2x = 2 ?

OpenStudy (anonymous):

That still wouldn't give any of the possible answers you have listed.

OpenStudy (ghazi):

3-2x=2......3-2=2x....now you can solve for x

OpenStudy (anonymous):

Oh, I see now.

OpenStudy (anonymous):

Just one sec.

OpenStudy (anonymous):

You have to use logarithms to solve this problem.

jimthompson5910 (jim_thompson5910):

everything is correct but the final answer is positive

OpenStudy (anonymous):

Whoops, Jim is right.

jimthompson5910 (jim_thompson5910):

\[\Large 3^{1-2x} = 2\] \[\Large 1-2x = \log_{3}(2)\] \[\Large -2x = \log_{3}(2) - 1\] \[\Large x = -\frac{\log_{3}(2) - 1}{2}\] \[\Large x \approx 0.18453512321428\] \[\Large x \approx 0.185\]

OpenStudy (anonymous):

How did you get that Jim I have no idea how to

jimthompson5910 (jim_thompson5910):

which step is throwing you off?

OpenStudy (anonymous):

Everything lol

jimthompson5910 (jim_thompson5910):

Step 1 ... Start with the original equation Step 2 ... Convert to logarithmic form Step 3 ... Subtract 1 from both sides. Step 4 ... Divide both sides by -2. Step 5 ... Use a calculator to approximate the right side Note: you type in "-(log(2)/log(3)-1)/(2)" without quotes Step 6 ... Round to 3 decimal places

jimthompson5910 (jim_thompson5910):

Let me know if that helps